# Neuromorphic Attitude Estimation and Control

### v1

@@ III-C Power usage analysis
The main benefits of having an end-to-end attitude SNN mainly derive from its combination with other autonomy functions such as computer vision on a single neuromorphic chip. Given the elementary nature of attitude estimation and control task, we do not expect any substantial performance or energy improvements for attitude estimation and control by itself. Still, we do think it is insightful to analyze the power usage of the current solution. The SNN in this research runs on a conventional microprocessor, as currently available neuromorphic chips (like Intel’s Loihi [28, 41] or SpiNNaker [42]) require supporting embedded systems that are too large for a 35-gram quadrotor or challenging to source. To explore potential power advantages, we performed some estimative calculations. Spike propagation through the network relies solely on additions rather than multiplications, allowing us to calculate the necessary operations based on addition alone. For the three-layer network used here, this would initially amount to approximately 42,500 additions per update. However, due to the 15-20% sparsity in neuron activations at each timestep, the actual required operations reduce to around 7,500 additions. At first glance, this still seems high compared to the traditional complementary filter and cascaded PID controller on the Crazyflie, which only requires about 28 additions per timestep. When including multiplications, however, the comparison changes: a 32-bit floating-point multiplication uses roughly 37 times more energy than a 32-bit integer addition [43]. Thus, the 52 multiplications in the conventional controller equate to an approximate cost of 2,000 additions per timestep. This suggests that, even with current hardware limitations, a neuromorphic algorithm like ours performs in the same energy order of magnitude as a PID-based controller.
Expanding this network to handle image data, as seen in other neuromorphic works (e.g. [29]), would create larger disparities due to the high multiplication demands in image processing tasks. These advantages align well with event-based control, which has demonstrated potential for drastic reductions in computational load (up to 80% for quadrotor attitude control [44]) by activating only when significant events occur.
@@ IV Conclusion
In this article, we have presented the first fully spiking attitude estimation and control pipeline for a quadrotor. We show that by using imitation learning, it is possible to train a fully end-to-end SNN to control a micro drone. We augmented training data to further enhance the performance, using in-flight data. The network was also taught to predict a $k$-step advance control action to mitigate delays that are inherent to the SNN. These methods led to significant reductions in RMSE relative to the target attitude and decreased oscillations, collectively enhancing the drone’s flight stability. Furthermore, our findings indicate that constraining parameters during training to function as integrators improves training precision and information integration. For RNNs these parameters would be the recurrent weights, and for SNNs the leak and threshold parameters. This novel approach avoids local minima during training and allows for faster convergence. By evaluating the system’s performance in real-world conditions and comparing it with traditional control methods, we have laid the groundwork for future developments in neuromorphic control strategies. Future research should aim to implement these algorithms on neuromorphic hardware, which could yield substantial gains in energy efficiency and reduced latency, potentially extending flight times and enabling neuromorphic UAVs in energy-constrained scenarios. By advancing these techniques, we envision the next generation of highly efficient, adaptive, and intelligent UAVs.
Acknowledgment
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-20-1-7044.
@@ Acknowledgment
All code necessary to 1) train the SNN, 2) convert and run the SNN on a Teensy 4.0, 3) integrate in the Crazyflie firmware and 4) perform the tests can be found in https://github.com/tudelft/neuromorphic_att_est_and_control. The data that was used for training can be found in https://data.4tu.nl/. The code and data will be made available upon publication.

### v2

@@ III-C Power usage analysis
The main benefits of having an end-to-end attitude SNN mainly derive from its combination with other autonomy functions such as computer vision on a single neuromorphic chip. Given the elementary nature of attitude estimation and control tasks, we do not expect any substantial performance or energy improvements for attitude estimation and control by itself.
Still, we do think it is insightful to analyze the power usage of the current solution. The SNN in this research runs on a conventional microprocessor, as currently available neuromorphic chips (like Intel’s Loihi [28, 42] or SpiNNaker [43]) require supporting embedded systems that are too large for a 35-gram quadrotor or challenging to source. To explore potential power advantages, we performed some estimative calculations. Spike propagation through the network relies solely on additions rather than multiplications, allowing us to calculate the necessary operations based on addition alone. For the three-layer network used here, this would initially amount to approximately 42,500 additions per update. However, due to the 15-20% sparsity in neuron activations at each timestep, the actual required operations reduce to around 7,500 additions. In contrast, the cascaded PID controller on the Crazyflie requires about 28 additions and 52 multiplications per timestep. Moreover, a straightforward complementary attitude estimation filter will have as most expensive operation a non-linear atan2 function that requires in the order of 15-30 multiplications. Since a 32-bit floating-point multiplication uses roughly 37 times more energy than a 32-bit integer addition [44], we can roughly equate the number of additions of a straightforward traditional pipeline with $\approx$ 3,000 additions. Hence, on a conventional microcontroller, the SNN performs in the same energy order of magnitude as a PID-based controller.
If small neuromorphic hardware becomes available that can natively support IMU readings, while implementing the SNN in hardware, energy consumption can be substantially reduced. Nonetheless, we maintain that the real gain would come when expanding this network to handle image data for instance, as seen in other neuromorphic works that show up to 100 $\times$ gains in efficiency (e.g. [29, 42]). This would create larger disparities due to the high multiplication demands in image processing tasks. Then, implementing all functionality in a single neuromorphic chip would make conventional companion computers obsolete, massively reducing energy consumption.
Finally, further benefits can be expected when moving to event-based control, which has demonstrated potential for drastic reductions in computational load (up to 80% for quadrotor attitude control [45]) by activating only when significant events occur. A drone in hover should only need to interfere and adapt its actuator commands when it starts to move, requiring no energy expenditure in between control events. Current microprocessors can not optimally benefit, because they still need to perform operations at a fixed frequency.
@@ IV Conclusion
In this article, we have presented the first fully spiking attitude estimation and control pipeline for a quadrotor. We show that by using imitation learning, it is possible to train a fully end-to-end SNN to control a micro drone. We augmented training data to further enhance the performance, using in-flight data. The network was also taught to predict a $k$-step advance control action to mitigate delays that are inherent to the SNN. These methods led to significant reductions in RMSE relative to the target attitude and decreased oscillations, collectively enhancing the drone’s flight stability. Furthermore, our findings indicate that constraining parameters during training to function as integrators improves training precision and information integration. For RNNs these parameters would be the recurrent weights, and for SNNs the leak and threshold parameters. This novel approach avoids local minima during training and allows for faster convergence. Next to that, our methods of implicitely learning integration and differentiation are not only applicable to attitude control for quadrotors, but apply to perception and control for robotics in general (e.g. using integration with rotary encoders or using differentiation to predict future states in model-based control). By evaluating the system’s performance in real-world conditions and comparing it with traditional control methods, we have laid the groundwork for future developments in neuromorphic control strategies. The importance of a working imitation learning pipeline, for instance, has been demonstrated in [46], where the authors show that bootstrapping a RL pipeline with imitation learning results in more reliable RL training while outperforming imitation learning only. Our methods can thus be used to improve RL for SNNs.
Future research should aim to implement these algorithms on neuromorphic hardware, which could yield substantial gains in energy efficiency and reduced latency, potentially extending flight times and enabling neuromorphic UAVs in energy-constrained scenarios. By advancing these techniques, we envision the next generation of highly efficient, adaptive, and intelligent UAVs.
@@
All code necessary to 1) train the SNN, 2) convert and run the SNN on a Teensy 4.0, 3) integrate in the Crazyflie firmware and 4) perform the tests can be found in https://github.com/tudelft/neuromorphic_att_est_and_control. The data that was used for training can be found here https://doi.org/10.4121/f474ef0a-6ef1-4ea1-a958-4827c4eadf60.